OpenStudy (anonymous):

solve equation: 38=26x-7x

OpenStudy (anonymous):

X=2

OpenStudy (anonymous):

howd u get that?

OpenStudy (anonymous):

26x-7x =19x so you have 19x=38 and to solve for x you do 38/19 which equals 2

OpenStudy (anonymous):

26x - 7x = 38 19x = 38 x = 38/19 x = 2